OpenStudy (anonymous):

lim tan(9/4pi-pisint/t) t->0 I will post equation in comments to make it easier to see

OpenStudy (anonymous):

\[limt \rightarrow \tan((9/4)\pi-(\pi \sin(t)/t)\]

zepdrix (zepdrix):

I can't understand your brackets. Lemme see if this is what you meant...

zepdrix (zepdrix):

\[\Large\bf\sf \lim_{t\to0}\tan\left(\frac{9\pi}{4}-\frac{\pi \sin t}{t}\right)\]

OpenStudy (anonymous):

ya

zepdrix (zepdrix):

Let's pass the limit into the tangent ( I think we're allowed to do that ) <.< \[\Large\bf\sf \tan\left(\lim_{t\to0} \frac{9\pi}{4}-\frac{\pi \sin t}{t}\right)\]And let's just focus on what's going on inside of the brackets for a moment,\[\Large\bf\sf \lim_{t\to0} \frac{9\pi}{4}-\frac{\pi \sin t}{t}\]

zepdrix (zepdrix):

Recall that for small t values, \(\Large\bf\sf \sin t\approx t\) Or you can just remember this identity if it makes more sense to you,\[\Large\bf\sf \lim_{x\to0}\frac{\sin x}{x}\quad=\quad 1\]

OpenStudy (anonymous):

ya i remember that

zepdrix (zepdrix):

So what do we get for this? :o\[\Large\bf\sf \lim_{t\to0} \frac{9\pi}{4}-\pi\color{royalblue}{\frac{ \sin t}{t}}\]

OpenStudy (anonymous):

1

zepdrix (zepdrix):

Not the whole thing, right? Just the sint/t\[\Large\bf\sf \lim_{t\to0} \frac{9\pi}{4}-\pi\color{royalblue}{\frac{ \sin t}{t}}\quad=\quad \frac{9\pi}{4}-\pi\color{royalblue}{(1)}\]

OpenStudy (anonymous):

ya how would you simplify

zepdrix (zepdrix):

\[\Large\bf\sf \tan\left(\frac{9\pi}{4}-\pi\right)\]Just need to get a common denominator. Or if you remember that tangent is periodic in pi, you can probably take advantage of that to skip a step.

OpenStudy (anonymous):

5pi/4?

zepdrix (zepdrix):

ya sounds good. Remember your coordinates on the unit circle? :)

OpenStudy (anonymous):

would the answer just be tan(5pi/4) or do i need more

OpenStudy (anonymous):

cause i don't know what tan(5pi/4) is

zepdrix (zepdrix):

You should know! :U

zepdrix (zepdrix):



zepdrix (zepdrix):

Something to remember: Tangent is always 1 or -1 at the pi/4 values. Is tangent positive or negative in the 3rd quadrant?

OpenStudy (anonymous):

positive

zepdrix (zepdrix):

Ok good, so we simply get a positive 1.\[\Large\bf\sf \tan(5\pi/4)=1\]

OpenStudy (anonymous):

ok theres one more really complicated one i have no clue how to solve

OpenStudy (anonymous):

\[4\pi \sec(3e^x+2\pi \tan(3\pi/secx)-3\]

OpenStudy (anonymous):

as x goes to 0

zepdrix (zepdrix):

You're missing a bracket, is it kind of like this? \[\Large\bf\sf \lim_{x\to0}4\pi \sec\left(3e^x+2\pi \tan\left(\frac{3\pi}{\sec x}\right)\right)-3\]

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

wait the -3 at the end is inside the last bracket

zepdrix (zepdrix):

\[\Large\bf\sf \lim_{x\to0}4\pi \sec\left(3e^x+2\pi \tan\left(\frac{3\pi}{\sec x}\right)-3\right)\]

OpenStudy (anonymous):

ya thats it

zepdrix (zepdrix):

This one is kind of nice. We can simply plug in x=0 without any problem.

zepdrix (zepdrix):

\[\Large\bf\sf \sec 0\quad=\quad \frac{1}{\cos 0}\quad=\quad ?\]

OpenStudy (anonymous):

1

zepdrix (zepdrix):

\[\Large\bf\sf \lim_{x\to0}4\pi \sec\left(3e^x+2\pi \tan\left(\frac{3\pi}{1}\right)-3\right)\]Ok good.

zepdrix (zepdrix):

Any other simplification that we can make? :o

OpenStudy (anonymous):

tan(3pi)=0?

OpenStudy (anonymous):

wait never mind

zepdrix (zepdrix):

Yah that sounds good. What can we do with the e^x part?

zepdrix (zepdrix):

\[\Large\bf\sf \lim_{x\to0}4\pi \sec\left(3e^x+2\pi(0)-3\right)\]

OpenStudy (anonymous):

would you plug 0 in the x

zepdrix (zepdrix):

ya

OpenStudy (anonymous):

so thats just 3

OpenStudy (anonymous):

so the inside is 0

OpenStudy (anonymous):

so the answer is 4pi?

zepdrix (zepdrix):

Mmmm ya I think that's right. I'll check it on wolfram a sec.

zepdrix (zepdrix):

yay 4pi \c:/ gj

OpenStudy (anonymous):

thanks!!!!!!